OpenStudy (anonymous):

Factor 25x2 + 10x + 1

OpenStudy (anonymous):

(5x +1)^2 = (5x+1)(5x+1) = 25x^2 + 5x + 5x +1 = 25x^2 + 10x +1

OpenStudy (anonymous):

A.(25x + 1)(x + 1) B.(5x + 1)(5x - 1) C.(5x + 1)² answer choices

OpenStudy (anonymous):

I gave it to you above.

jhonyy9 (jhonyy9):

c. is correct